# Constrained Bandwidth Observation Sharing for Multi-Robot Navigation in Dynamic Environments via Intelligent Knapsack

### v1

@@ Abstract
Multi-robot navigation is increasingly crucial in various domains, including disaster response, autonomous vehicles, and warehouse and manufacturing automation. Robot teams often must operate in highly dynamic environments and under strict bandwidth constraints imposed by communication infrastructure, rendering effective observation sharing within the system a challenging problem. This paper presents a novel optimal communication scheme, Intelligent Knapsack (iKnap), for multi-robot navigation in dynamic environments under bandwidth constraints. We model multi-robot communication as belief propagation in a graph of inferential agents. We then formulate the combinatorial optimization for observation sharing as a 0/1 knapsack problem, where each potential pairwise communication between robots is assigned a decision-making utility to be weighed against its bandwidth cost, and the system has some cumulative bandwidth limit. Compared to state-of-the-art broadcast-based optimal communication schemes, iKnap yields significant improvements in navigation performance with respect to scenario complexity while maintaining a similar runtime. Furthermore, iKnap utilizes allocated bandwidth and observational resources more efficiently than existing approaches, especially in very low-resource and high-uncertainty settings. Based on these results, we claim that the proposed method enables more robust collaboration for multi-robot teams in real-world navigation problems.
@@ I Introduction
Most recently, Marcotte et al. [18] proposed a method for optimizing multi-robot communication under bandwidth constraints using a bandit-based combinatorial optimization technique to select observations for inclusion in messages. However, the formulation relies on a broadcast-based communication scheme where messages are shared with all agents simultaneously. In practice, the utility of a particular piece of information for decision-making is often highly subjective to each individual agent. For example, the observation of some obstacle might only be useful to agents whose planned paths potentially run into that obstacle, and irrelevant to those far away. In that case, by sending messages only to those who need them, we can save communication bandwidth as well as computation power without sacrificing performance. Thus, we are motivated to formalize the above intuition into a pairwise communication scheme.
@@ III-C Combinatorial Optimization Problem
This problem takes the form of a 0/1 knapsack problem, which is solvable in pseudo-polynomial time via dynamic programming. In particular, the algorithm finds a subset of $N$ items, each with some weight and value, to place in a knapsack of weight capacity $B$ such that the cumulative value is maximized in $O(NB)$ time. In our case, $N$ is $O(n^{2})$, yielding a time complexity of $O(n^{2}B)$. For a depiction of the overall observation sharing process, see Algorithm 1. For an example of the observation uploading process and a comparison of pairwise and broadcast-based communication schemes, see 2.
@@ IV Experimental Results
In this section, we evaluate the proposed method under various simulated conditions and compare its performance to the current state of the art.
@@ IV-A Simulation Setup
We design a randomized Python simulation environment with an open field containing $n$ robots, $m$ dynamic obstacles as subjects of observation, and $w$ walls as static obstacles to maneuver around. Robots and subjects follow 2D double-integrator dynamics from starting locations to goal locations. Robots obtain noisy observations of visible subjects around them and update their beliefs accordingly. We also implement a centralized communication infrastructure that collects uploaded observations and beliefs from each agent, runs iKnap to compute optimal communications, and executes these communications by sending observations to the relevant agents at some fixed frequency.
@@ IV-B Perception Design
At each timestep, each agent obtains a noisy observation of each visible subject’s state, i.e. position and velocity. We define a subject to be visible to an agent if no wall exists between the subject and the agent, which reflects real-world occlusion scenarios. Then, between agent $i$ and subject $k$, the observation $\boldsymbol{\mu}_{ik}$ is corrupted with Gaussian noise
@@
for tuning parameter $\alpha$, i.e. observations are less accurate the further away an agent is from a subject. Each time agent $i$ receives a new observation of subject $k$’s state, it updates its beliefs $\boldsymbol{\Phi}_{ik}$ via Bayesian inference with Gaussian prior and likelihood and assuming constant velocity. An agent’s confidence in its observation is proportional to the perceived distance between itself and the subject.
IV-C Utility Heuristic Design
The utility heuristic must provide an estimate of how useful an observation from agent $a$ regarding subject $h$ would be to agent $b$. Intuitively, we must take into account
  • the improvement in $b$’s belief about $h$ would improve as a result of $a$’s observation, and
  • the relevance of knowledge about $h$’s state in $j$’s navigation task.
For the former criterion, we choose to measure the Kullback–Leibler divergence between the prior and posterior distributions of $\boldsymbol{\Phi}_{bh}$ for position and velocity, using $a$’s observation as the likelihood distribution; we denote this value $\kappa_{abh}$. For the latter criterion, we choose to forward-simulate the trajectories of $b$ and $h$, assuming a constant-velocity trajectory for $h$ based on $a$’s beliefs, and then measure the inverse-squared minimum distance between $b$ and $h$ across a horizon of $t_{H}$ seconds in the navigation task; we denote this value $\tau_{abh}$. Then, the utility for potential communication $C_{i}$ associated with $(a,b,h)$ is given by
@@ IV-C Utility Heuristic Design
for tuning parameters $p_{1}$ and $p_{2}$ to normalize and determine the relative importance of each metric.
IV-D Planner and Controller Design
Each agent plans a straight-line path from its start location to its goal location that wraps around walls where necessary. By default, each agent has some target speed along its trajectory. However, each agent runs a trajectory predictor assuming constant velocity based on its beliefs about the state of nearby subjects, and if it perceives an impending collision with a subject, it sets its target speed to 00 at the location of the collision. The distance-to-collision computation returns a probability distribution propogated from the uncertainty in the subject’s state, and the agent subtracts one standard deviation from the mean distance in order to ensure safety. The agent uses PD control to compute the corresponding acceleration at each timestep. Notice that this implementation rewards an agent’s greater confidence in its observations of subjects, as this directly improves navigation outcomes.
For each subject, we generate a randomized cubic spline trajectories from its start location to its goal location. We set some random target speed for the subject along its trajectory, and the subject uses PD control to compute the corresponding acceleration at each timestep.
IV-E Baseline and Evaluation Metrics
We compare the proposed framework’s performance to that of OCBC, a state-of-the-art observation sharing method proposed in [18]. Most notably, OCBC operates under a broadcast-based communication scheme, as opposed to a pairwise scheme. Additionally, OCBC leverages a bandit-based combinatorial optimization approach for selecting observations to share, as opposed to a knapsack approach. Consequently, OCBC is unable to account for variations in bandwidth required for each potential communication.
We measure algorithm performance via runtime where necessary, and we measure navigation performance via makespan, which is justified from the agent controller design presented in Section IV-D. All experiments were run on a desktop computer with an AMD Ryzen 5 2600X CPU and 16GB RAM. In the following sections, we first evaluate the capability of iKnap, compared to both OCBC and a no-communication control condition, and then we perform ablation studies comparing iKnap to OCBC.
(a) Agent/subject count $n+m$
@@ IV-E Baseline and Evaluation Metrics
(c) Wall count $w$
(d) Optimizer runtime (s)
Figure 3: Capability studies for iKnap (blue), a no-communication control condition (orange), and OCBC (green). We measure makespan (s) vs. agent/subject count $n+m$, noise parameter $\alpha$, and wall count $w$. We also measure optimizer runtime (s) vs. agent/subject count $n+m$. Each configuration’s measurements are averaged across $100$ randomized trials.
@@
(b) Bandwidth range
(c) Max. bandwidth proportion
(d) Prediction horizon (s)
Figure 4: Ablation studies for iKnap (blue) and OCBC (green). We measure makespan (s) vs. communication frequency (hz), range of randomly generated pairwise bandwidth values, bandwidth limit ratio $B/n$, and agent trajectory prediction horizon (s). Each configuration’s measurements are averaged across $100$ randomized trials.
IV-F Capability Studies
To assess the capability of iKnap, we measure the navigation performance with respect to the size of the communication problem, environment complexity, and observation noise. Thus, we measure makespan with respect to the combined number of agents and subjects in equal number, the number of walls in the environment, and the amount of agent sensor noise, respectively. We also measure the runtime of the communication scheme with respect to the agent and subject count. We compare iKnap’s performance to that of OCBC, as well as a no-communication baseline to exemplify the effectiveness of observation sharing between agents. By default, we configure $n=5$, $m=5$, $w=10$, and $\alpha=0.01$. For all other default values, see Section IV-G. For each parameter configuration, we simulate $100$ trials with randomized initial conditions and compute the average value.
Figure 3 displays the results of the capability experiments. These results imply that iKnap’s navigation performance scales better than OCBC and the no-communication control group with respect to scenario complexity (Fig. 3a, Fig. 3c) and under significant observational constraints (Fig. 3b). Moreover, Figure 3d displays the optimizer runtimes for iKnap and OCBC with respect to scalability. The two algorithms exhibit largely similar runtime, and even in scenarios involving $20$ robots and $20$ dynamic obstacles as subjects, iKnap poses only a $10\%$ increase in runtime over OCBC, while providing a consistent increase of $20\%-35\%$ in navigation performance across the other measured variables.
IV-G Ablation Studies
To perform ablation studies for iKnap, we measure the navigation performance with respect to communication frequency, bandwidth allowance, bandwidth variability, and utility heuristic performance. Thus, we measure makespan with respect to the frequency at which the optimal communication problem is run and its results executed, the ratio of the bandwidth limit to the number of agents, the range of randomly generated bandwidth costs associated with communications between each pair of agents, and the trajectory prediction horizon for our chosen utility heuristic. We compare iKnap’s performance to that of OCBC. By default, we configure a communication frequency of $1$ hz, a bandwidth limit ratio of $5$ units per agent, uniformly random pairwise bandwidths from $1$ to $10$ units (yielding a range of $9$ units), and a prediction horizon of $5$ seconds. For each parameter configuration, we simulate $100$ trials with randomized initial conditions and compute the average value.
Figure 4 displays the results of the ablation experiments. These results imply that iKnap utilizes allocated communication resources (Fig. 4a-c) and observational resources (Fig. 4d) more effectively than OCBC. Moreover, iKnap excels in heavily constrained scenarios.

### v2

@@ Abstract
Multi-robot navigation is increasingly crucial in various domains, including disaster response, autonomous vehicles, and warehouse and manufacturing automation. Robot teams often must operate in highly dynamic environments and under strict bandwidth constraints imposed by communication infrastructure, rendering effective observation sharing within the system a challenging problem. This paper presents a novel optimal communication scheme, Intelligent Knapsack (iKnap), for multi-robot navigation in dynamic environments under bandwidth constraints. We model multi-robot communication as belief propagation in a graph of inferential agents. We then formulate the combinatorial optimization for observation sharing as a 0/1 knapsack problem, where each potential pairwise communication between robots is assigned a decision-making utility to be weighed against its bandwidth cost, and the system has some cumulative bandwidth limit. We evaluate our approach in a simulated robotic warehouse with human workers using ROS2 and the Open Robotics Middleware Framework. Compared to state-of-the-art broadcast-based optimal communication schemes, iKnap yields significant improvements in navigation performance with respect to scenario complexity while maintaining a similar runtime. Furthermore, iKnap utilizes allocated bandwidth and observational resources more efficiently than existing approaches, especially in very low-resource and high-uncertainty settings. Based on these results, we claim that the proposed method enables more robust collaboration for multi-robot teams in real-world navigation problems.
@@ I Introduction
Most recently, Marcotte et al. [18] proposed a method for optimizing multi-robot communication under bandwidth constraints using a bandit-based combinatorial optimization technique to select observations for inclusion in messages. Their approach, OCBC, represents the state of the art due to its efficient consideration of observation utility. However, the formulation relies on a broadcast-based communication scheme where messages are shared with all agents simultaneously. In practice, the utility of a particular piece of information for decision-making is often highly subjective to each individual agent. For example, the observation of some obstacle might only be useful to agents whose planned paths potentially run into that obstacle, and irrelevant to those far away. In that case, by sending messages only to those who need them, we can save communication bandwidth as well as computation power without sacrificing performance. Thus, we are motivated to formalize the above intuition into a pairwise communication scheme.
@@ III-B Pairwise Communication Scheme
One key assumption we make when formulating this communication scheme is that the bandwidth cost of a broadcast can be approximated as the sum of the bandwidth costs of each individual pairwise communications. This assumption is well-supported by empirical studies in multi-robot communications. For example, broadcast transmissions require power proportional to maximum sender-receiver distance, while directed communications optimize based on specific distances [21]. Moreover, broadcasts suffer increased packet loss in larger teams [22], with pairwise schemes achieving significant bandwidth reduction compared to broadcast approaches [4]. These findings from real-world deployments validate our modeling approach.
@@ III-C Combinatorial Optimization Problem
This problem takes the form of a 0/1 knapsack problem, which is solvable in pseudo-polynomial time via dynamic programming. In particular, the algorithm finds a subset of $N$ items, each with some weight and value, to place in a knapsack of weight capacity $B$ such that the cumulative value is maximized in $O(NB)$ time. In our case, $N$ is $O(n^{2})$, yielding a time complexity of $O(n^{2}B)$. For a depiction of the overall observation sharing process, see Algorithm 1. For an example of the observation uploading process and a comparison of pairwise and broadcast-based communication schemes, see Figure 2.
@@ IV Experimental Results
Figure 3: Example warehouse environment generated with the Open-RMF Traffic Editor tool.
@@ IV-A Simulation Setup
We design a high-fidelity multi-robot warehouse simulation environment using the Robot Operating System (ROS 2) [23] and the Open Robotics Middleware Framework (RMF) [24]. We choose RMF for its comprehensive capabilities in simulating the full multi-robot navigation stack. The environment consists of a warehouse with $n$ robots, $m$ dynamic obstacles (simulated human workers), and various static obstacles (walls, shelves) to navigate around. The simulation leverages RMF’s traffic_editor package to generate randomized warehouse environments with realistic layouts. Robots follow paths computed by a grid-based A* planner and use the RMF slotcar plugin for dynamics and control. This plugin provides a lightweight but realistic model of differential-drive robot kinematics. Human workers are simulated as dynamic obstacles using RMF’s CrowdSim package, which provides realistic human movement patterns in warehouse environments. The humans follow randomized paths between different locations in the warehouse, creating dynamic navigation challenges for the robots. All simulations are run using $\Delta t=0.01$ s. For an example warehouse environment map generated by RMF, see Figure 3.
@@ IV-B Perception Design
At each timestep, each robot obtains a noisy observation of each visible dynamic obstacle’s state (position and velocity). An obstacle is considered visible to a robot if no walls or shelves exist between them, accurately modeling real-world occlusion scenarios. The observation noise is modeled as a Gaussian distribution with standard deviation
@@
for tuning parameter $\alpha$, reflecting the realistic property that observations become less accurate with increasing distance.
Each robot maintains a belief state about each dynamic obstacle using Bayesian inference with Gaussian prior and likelihood, assuming constant velocity dynamics. When a robot receives a new observation (either from its own sensors or shared by another robot), it updates its belief accordingly.
IV-C Communication Infrastructure
We implement a centralized communication infrastructure that collects observations from all robots, determines optimal sharing strategies, and distributes selected observations. The infrastructure models realistic communication constraints through several key factors. The bandwidth required for communication between robots $a$ and $b$ is modeled as a function of their Euclidean distance, occlusions, and message size:
$\beta_{ab}=\beta_{base}+\beta_{dist}\cdot||x_{a}-x_{b}||+\beta_{occl}\cdot\mathbb{I}_{occl}(a,b)$ (4)
where $\beta_{base}$ is a base cost, $\beta_{dist}$ is a distance-dependent factor, and $\beta_{occl}$ is an additional cost applied when occlusions exist between robots. The system also models communication latency as a configurable delay between when a message is sent and when it is received, simulating real-world network conditions. Additionally, a global constraint on the total bandwidth available per communication cycle is implemented to reflect realistic limitations.
IV-D Utility Heuristic Design
The utility heuristic provides an estimate of how useful an observation from robot $a$ regarding subject $h$ would be to robot $b$. We consider two primary factors in this calculation. First, we measure the improvement in $b$’s belief about $h$ that would result from $a$’s observation, using the Kullback–Leibler divergence between prior and posterior distributions of $\boldsymbol{\Phi}_{bh}$, denoted as $\kappa_{abh}$. Second, we assess the relevance of knowledge about $h$’s state to $b$’s navigation task by forward-simulating trajectories and computing the inverse-squared minimum distance between $b$ and $h$ across a prediction horizon of $t_{H}$ seconds, denoted as $\tau_{abh}$. The utility for potential communication $C_{i}$ associated with $(a,b,h)$ is given by
@@ IV-D Utility Heuristic Design
for tuning parameters $p_{1}$ and $p_{2}$ that normalize and determine the relative importance of each metric. Intuitively, $\kappa_{abh}$ measures the improvement in $b$’s knowledge, and $\tau_{abh}$ measures the applicability of $b$’s knowledge. Taken together, these two factors form a reasonable heuristic for the true “utility” of a given communication. We note that there are many other ways to choose the utility heuristic, e.g. via learning-based approaches.
IV-E Planner and Controller Design
Each robot plans a path from its start location to its goal location using a grid-based A* algorithm that accounts for static obstacles. The robots follow these paths using the RMF slotcar plugin, which implements an adapted model of differential-drive robot dynamics. For collision avoidance with dynamic obstacles, each robot leverages its belief about nearby obstacles to guide naive trajectory prediction with a constant-velocity assumption. If a collision is predicted, the robot decelerates linearly or stops until its confidence threshold path is clear is satisfied. The trajectory predictor is implemented with a variable prediction horizon parameter. This implementation rewards robots having greater confidence in their observations, as this directly improves navigation outcomes by reducing unnecessary slowdowns.
IV-F Dynamic Obstacles Design
To simulate human workers in the warehouse environment, we leverage RMF’s CrowdSim package which integrates the Menge library for realistic human movement. Dynamic obstacles are modeled as circular agents with configurable radii (0.25m) navigating through a graph-based global plath planner and Optimal Reciprocal Collision Avoidance (ORCA) [25] for local collision avoidance. Human movements follow a Finite State Machine where agents transition between goal states based on proximity or time conditions. Each human has individualized parameters including maximum acceleration (5 m/s2), maximum angular velocity (360 deg/s), preferred speed (1.5 m/s), and maximum speed (2 m/s). This implementation creates non-deterministic human movements that respond dynamically to both the environment and other agents, providing realistic warehouse interaction scenarios that challenge the robots’ navigation capabilities.
IV-G Baselines and Evaluation Metrics
We compare the iKnap framework to four baseline approaches. The No Communication baseline has robots rely solely on their own observations without sharing information. The Random Broadcast approach has robots randomly decide whether to broadcast their observations to all other robots with a fixed probability. In the Random Pairwise method, robots randomly decide whether to share observations with each specific robot with a fixed probability. Finally, OCBC is a state-of-the-art observation sharing method proposed in [18] that uses a broadcast-based communication scheme with a bandit-based combinatorial optimization approach for utility-based observation selection.
We measure algorithm performance via two primary metrics: makespan, which is the time until the last robot reaches its goal and directly reflects the overall efficiency of the multi-robot system; and optimizer runtime, which is the computational time required by the communication selection algorithm and is important for real-time applications.
All experiments were run on a Linux desktop computer with an Intel Core i7-12700KF CPU and 16GB DDR5 RAM. In the following sections, we first evaluate the capability of iKnap compared to all baselines, and then perform a robustness analysis comparing primarily to OCBC.
(a) Robot/obstacle count $n+m$
@@ IV-G Baselines and Evaluation Metrics
(c) Communication latency
(d) Wall count $w$
Figure 4: Capability studies comparing makespan (s) for iKnap (blue), OCBC (green), random pairwise (red), random broadcast (orange), and no communication (purple) baselines. We measure performance vs. (a) robot/obstacle count $n+m$, (b) sensor noise level, (c) communication latency, and (d) environment complexity (number of walls). Each configuration’s measurements are averaged across multiple randomized trials.
IV-H Capability Studies
To assess the capability of iKnap, we measure navigation performance with respect to the size of the communication problem, environment complexity, observation noise, and communication latency. We compare our approach to all baselines across these dimensions. For each parameter configuration, we average performance across $20$ randomly generated warehouse environment configurations. When altering one parameter, we hold all other parameters constant at half of their maximum tested value.
Figure 4 displays the results of the capability experiments. These results demonstrate that iKnap consistently outperforms all baselines across all tested conditions. As expected, the no-communication baseline performs worst, while the random communication approaches show modest improvements. OCBC performs significantly better than the random approaches but is consistently outperformed by iKnap. Notably, randomized pairwise communication poses a consistent advantage over randomized broadcast-based communication, highlighting the fundamental differences between these formulations.
Our observations specifically indicate that our approach offers significant advantages in several key areas. In terms of scalability (Fig. 4a), as the number of agents and obstacles increases, all methods show increased makespan due to greater congestion and coordination challenges. However, iKnap scales more gracefully, maintaining a 1̃0% advantage over OCBC even at higher agent counts. Regarding robustness to noise (Fig. 4b), with increasing sensor noise, all methods degrade in performance. Our approach shows greater robustness, with the performance gap between iKnap and OCBC widening as noise increases, demonstrating the value of selective information sharing under uncertainty. For communication latency (Fig. 4c), higher latency negatively impacts all communication-based approaches. Our approach maintains better performance across all latency levels, suggesting more effective prioritization of time-sensitive information. In terms of environmental complexity (Fig. 4d), as the number of walls increases, creating more complex navigation scenarios with occlusions, iKnap maintains its advantage, showing 1̃5% lower makespan compared to OCBC.
Figure 6 shows the optimizer runtime for both iKnap and OCBC. Despite the theoretical complexity of the knapsack problem, our implementation exhibits runtime comparable to OCBC, with only a modest increase as the number of agents grows. Even in scenarios with 10 robots and 10 obstacles, the runtime remains well under 0.4 seconds, making it suitable for real-time applications.
Figure 5: Optimizer runtime (s) vs. agent/obstacle count for iKnap (blue) and OCBC (green).
@@ IV-H Capability Studies
(b) Bandwidth limit
(c) Prediction horizon (s)
Figure 6: Robustness analysis comparing makespan (s) for iKnap (blue), OCBC (green), random pairwise (red), and random broadcast (purple) approaches. We measure performance vs. (a) communication frequency, (b) bandwidth limit, and (c) trajectory prediction horizon. Each configuration’s measurements are averaged across multiple randomized trials.
IV-I Robustness Analysis
To analyze the robustness of iKnap, we measure navigation performance with respect to communication frequency, bandwidth allowance, and utility heuristic parameters. We compare primarily to OCBC, with the random approaches included for context.
Figure 6 displays the results of the robustness experiments. These results show several significant trends in system performance. Regarding communication frequency (Fig. 6a), higher frequency improves performance for all approaches by allowing more timely information sharing. Our approach shows greater improvement with increased frequency, suggesting iKnap selects more valuable information at each communication cycle. For bandwidth limits (Fig. 6b), as the limit increases, all approaches improve, but with diminishing returns. Our approach shows superior performance across all bandwidth limits, with the advantage being most pronounced at lower limits where selective sharing is most critical. In terms of prediction horizon (Fig. 6c), the horizon affects how far ahead robots consider potential interactions with dynamic obstacles. Both iKnap and OCBC show an optimal performance at a medium horizon (around 7 s), with performance degrading at very short horizons annd flattening out at very long horizons. This suggests that very short horizons provide insufficient lookahead, while very long horizons introduce too much uncertainty in predictions.
These robustness studies demonstrate that iKnap utilizes communication resources more effectively than OCBC and the random baselines. The advantage is particularly significant in resource-constrained scenarios (low communication frequency or bandwidth), highlighting the value of the pairwise communication paradigm and knapsack optimization for real-world multi-robot systems where communication resources are often limited.